OpenStudy (anonymous):

Can you tell me how to calculate this -9 mod26 i know the answer is 17 but have no idea why .

OpenStudy (jdoe0001):

what's the "mod"?

OpenStudy (anonymous):

modulo , from modulo arithmetic

OpenStudy (anonymous):

mainly used in cryptography

OpenStudy (anonymous):

Ok i figure it out. If anybody is interested, when dealing with negative number, we add the modulo until it becomes a positive integer. so -9 + 26 = 17 (mod 26). eg. -44 mod (26) will be -44+26 =-18 ; -18 + 26 = 8. so -44 mod 26 is 8